OpenStudy (anonymous):

Solve 3x2 = -30x + 24 PLEASE HELP ME! I didn't get the answer that's in the answer choices! these are the answer choices: quantity of negative 10 plus or minus square root of 33 all over 2 negative 5 plus or minus 2 square root of 33 negative 10 plus or minus square root of 33 negative 5 plus or minus square root of 33 and I got: -30-+sqr-1188/6

OpenStudy (anonymous):

\[3x^2 = -30x + 24\] \[3x^2+30x-24=0\] divide by 3 and get \[x^2+10x-8=0\]

OpenStudy (anonymous):

whaaaaat....?

OpenStudy (anonymous):

if you want to use the quadratic formula, set it equal to zero and solve

OpenStudy (anonymous):

I did, and I got the equation on top but I don't know what it is.

OpenStudy (anonymous):

if you start with \[3x^2 = -30x + 24\] then add \(30x\) and subtract \(24\) to get \[3x^2+30x-24=0\]

OpenStudy (anonymous):

then divide all by 3 because you don't need all those common factors

OpenStudy (anonymous):

\[x^2+10x-8=0\] makes the quadratic formula a lot easier to use

OpenStudy (anonymous):

okay then we get x^2+10x-8=0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now i solved and got -10-+sqr68/2

OpenStudy (anonymous):

in fact since \(10\) is even, it is a lot easier to complete the square than to use the quadratic formula

OpenStudy (anonymous):

\[x^2+10x=8\] \[(x+5)^2=8+25=33\] \[x+5=\pm\sqrt{33}\] and so \[x=-5\pm\sqrt{33}\]

OpenStudy (anonymous):

im lost..how did you get the second one?

OpenStudy (anonymous):

HELP!